OpenStudy (fernandez555):

(01.05) What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)? y = −3x − 2 y = −3x + 2 y = 3x + 2 y = 3x − 2

OpenStudy (anonymous):

first find the slope y2-y1/x2-x1 then use tangent line equation of y-y1 = m(x-x1) then put it into y=mx+b m is the slope

OpenStudy (the_fizicx99):

y2-y1/x2-x1 Plug in. After that you should get your slope (M) then do y-y1=m(x-x1) and that should give your y=mx+b

OpenStudy (anonymous):

y=3x-2

OpenStudy (anonymous):